OpenStudy (anonymous):

A plane travels at a speed of 180 mph in still air. Flying with a tailwind, the plane is clocked over a distance of 725 miles. Flying against a headwind, it takes 2 hours longer to complete the return trip. What was the wind velocity? (Round your answer to the nearest tenth.)

OpenStudy (misty1212):

you are endless i always get these backwards let me try and see if what i get makes sense

OpenStudy (anonymous):

I only have 2 or 3 more!

OpenStudy (anonymous):

Hopefully they are easy! :)

OpenStudy (misty1212):

i solved \[\frac{725}{180+x}=\frac{725}{180-x}+2\] but i got an answer that made no sense

OpenStudy (misty1212):

let me think about why it is wrong

OpenStudy (anonymous):

I got 18.97. I don't think mine is right

OpenStudy (misty1212):

with headwind rate is \(180+x\) so the time is \[\frac{725}{180+x}\]with the head wind

OpenStudy (anonymous):

Right I did that part

OpenStudy (misty1212):

maybe that was the tailwind in any case that was with the wind not against is

OpenStudy (misty1212):

against the wind the rate is \(180-x\) so the time is \(\frac{725}{180-x}\)

OpenStudy (misty1212):

ooh i see how i screwed up

OpenStudy (anonymous):

Yea and I subtracted them and set them equal to 2.

OpenStudy (misty1212):

i did this \[\frac{725}{180+x}=\frac{725}{180-x}+2\] but it should have been this \[\frac{725}{180+x}=\frac{725}{180-x}-2\]

OpenStudy (misty1212):

and got \(x=42.23\) which makes a lot more sense

OpenStudy (anonymous):

Yes it does!

OpenStudy (misty1212):

your math teacher must hate you

OpenStudy (anonymous):

I am taking prep for college math. I finished 3 semesters of work in 2 semesters but I had to get a little extension on the last part. She actually loved me because I tried!

OpenStudy (misty1212):

i am sure i just meant giving all these annoying word problems

OpenStudy (anonymous):

Yeah. Well it is online so she has no choice haha!

OpenStudy (anonymous):

I have been doing this math for 6 hours now. I am ready to be done!